OpenStudy (anonymous):

. A bag contains 7 red marbles, 2 blue marbles, and 1 green marble. What is the probability of choosing a marble that is not blue?

OpenStudy (anonymous):

the probability of choosing a marble that is not blue is 8/10 or 5/5 because there are 10 total marbles in the bag and 8 non blue marbles. you divide the total number of marbles by the amount of marbles that fit your criteria

OpenStudy (anonymous):

You mean 4/5...but everything else makes sense.

OpenStudy (raden):

or use the complement event : P(not blue) = 1 - P(blue) = 1 - 2/10 = 1 - 1/5 = 4/5